OpenStudy (anonymous):

find the slope and y intercept of the line in the given equation. y=6/7x-4 show work

OpenStudy (anonymous):

Do you have a graphing calculator handy?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Ah darn :/ I'm stuck on something like this myself. Lol :p

OpenStudy (anonymous):

slope =6/7 and y intercept is -4 From y=mx+b Where m is the slope and b is y-intercept

OpenStudy (anonymous):

slope is 6/7 and y intercept is -4., This is slope intercept for so, it's right in face!!!

OpenStudy (anonymous):

thank you. everyone on this site has been very helpful :)

OpenStudy (anonymous):

You're welcome!